OpenStudy (anonymous):

FINAL IN THE MORNING: the great circle of the sphere has a circumference of 24 (pic soon to be drawn in comments) what is the surface area

OpenStudy (anonymous):



OpenStudy (anonymous):

@FutureMathProfessor

OpenStudy (anonymous):

help me please @FutureMathProfessor

OpenStudy (anonymous):

@Goldpheonix

OpenStudy (anonymous):

Isn't this the same problem we helped you on earlier?

OpenStudy (anonymous):

yes I don't know how but it got deleted

OpenStudy (anonymous):

are you going to help me? @FutureMathProfessor

OpenStudy (anonymous):

Ok so you have to find R from the equation of circumfrence and then plug that into the equation of surface area to find the surface area

OpenStudy (anonymous):

how do I find r

OpenStudy (anonymous):

@oldrin.bataku are you good at these

OpenStudy (anonymous):

@FutureMathProfessor excuse me..

OpenStudy (anonymous):

@mathstudent55 can you please help me

OpenStudy (anonymous):

If the great circle has a circumference of \(24\), can you find its radius? This is also the radius of our sphere.

OpenStudy (anonymous):

so 24 is the radius

OpenStudy (anonymous):

no

OpenStudy (anonymous):

how do I find the radius

OpenStudy (anonymous):

\(C=2\pi r\) relates the circumference to the radius. If our circumference is \(24\), we have \(24=2\pi r\). Can you solve for \(r\)?

OpenStudy (anonymous):

ok so the circumference is 24 so would the radius be 12

OpenStudy (anonymous):

no

OpenStudy (anonymous):

not to be rude, but im going to need a way better answer than "no" @FutureMathProfessor

OpenStudy (anonymous):

You need to manipulate the circumfrence equation to solve for R.

OpenStudy (anonymous):

but im on here because I don't know how to do that...and your not telling me how to set it up @FutureMathProfessor

OpenStudy (anonymous):

@oldrin.bataku how d I find r

OpenStudy (anonymous):

@.Sam. any ideas?

OpenStudy (anonymous):

$$24=2\pi r\\12=\pi r\text{ after dividing both sides by }2\\\frac{12}\pi=r\text{ after dividing both sides by }\pi$$Now that we hav e \(r\), our surface area is given by \(A=4\pi r^2\).

OpenStudy (anonymous):

ok I still don't see the radius..

OpenStudy (anonymous):

\[A=4\Pi(\frac{ 12 }{ \Pi })^{2}\]

OpenStudy (anonymous):

still not seeing how I would get the radius

OpenStudy (anonymous):

any ideas @Tigger25

OpenStudy (anonymous):

Listed below is the equation for the circumference of a circle:\[C=2(\Pi)R\] Since we know C, we can plug it in and solve for R. \[24=2(\Pi)R\] Let's manipulate the equation to solve for R.\[R=\frac{ 24 }{ 2*\Pi }\] Use that to find R, then we will plug R into our Sphere Surface Area Equation, listed below:\[SA=4(\Pi)R^2\] Since we have our R now, we can just plug that into our new equation to find SA (Surface Area) and then we will have our answer!

OpenStudy (anonymous):

The radius is \[12/ \pi\] or 3.82 Then you substitute that in to the SA formula \[4 \pi r^2\] \[4\pi (3.82)^2\] and get 183.35

OpenStudy (anonymous):

@ anyone else please tell me if I made a mistake ^^

OpenStudy (anonymous):

I got two different answers and two different calculators...the format looks correct but im still not getting the answer

OpenStudy (anonymous):

mt teacher came out wit 576pie

OpenStudy (anonymous):

@Tigger25 ^

OpenStudy (anonymous):

oh wow all this time the radius was 12 lol @FutureMathProfessor I got the answer

OpenStudy (anonymous):

The radius is not 12!

OpenStudy (anonymous):

Can I have a medal? :D

OpenStudy (anonymous):

the radius was 12 that's how I got the CORRECT answer my teacher gave us..

OpenStudy (anonymous):

what are themedals for though do yal get paid to answer questions?

OpenStudy (anonymous):

The radius is 12/pi.............